OpenStudy (anonymous):

r = 1 - cos(theta) 0 < theta < 2pi find where the tangents are horizontal and vertical

OpenStudy (anonymous):

Find the derivative and set it equal to zero for horizontal.

OpenStudy (anonymous):

i get \[\frac{dy}{d \theta} = \sin^2\theta - \cos^2\theta + \cos \theta\]

OpenStudy (anonymous):

so its zero when theta is 0 , pi and 2pi right??

OpenStudy (anonymous):

\[r'=\sin(\Theta)\]

OpenStudy (anonymous):

WAIT! Are you in polar mode?

OpenStudy (anonymous):

You dont have to use \[x = rcos(\theta) \space and \space y= rsin(\theta)\] ??? and then plug in the r you have??

OpenStudy (anonymous):

Yes polar mode

OpenStudy (anonymous):

Yes you do - Ignore my responses. >.<

OpenStudy (anonymous):

Ok i did that and got the derivative that i typed above